40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Jan 24 13:55:04 2018, Kira, Summary, PEM, Seismometer can insulation test 6x
    Reply  Thu Jan 25 12:41:18 2018, Kira, Update, PEM, Seismometer can insulation test 
       Reply  Thu Jan 25 14:58:20 2018, Kira, Update, PEM, Seismometer can insulation test 
          Reply  Thu Jan 25 16:39:09 2018, Kira, Update, PEM, Seismometer can insulation test HeaterTest.png
             Reply  Thu Feb 1 16:09:13 2018, Kira, Update, PEM, Seismometer can insulation test cooling_fit.png
                Reply  Mon Feb 5 12:12:41 2018, Kira, Update, PEM, Seismometer can insulation test temp_data.py
                   Reply  Wed Feb 7 12:35:56 2018, Kira, Update, PEM, Seismometer can insulation test fit_1.png
                      Reply  Wed Feb 7 17:01:25 2018, gautam, Update, PEM, PID test plan ADC_EX.pngDAC_EX.pngIMG_20180207_170628.jpgdac_adc.jpgIMG_20180207_165833.jpg
                         Reply  Wed Feb 7 19:14:19 2018, rana, Update, PEM, PID test plan 
                            Reply  Thu Feb 8 12:24:37 2018, Kira, Update, PEM, PID test plan 
                               Reply  Thu Feb 8 17:32:44 2018, Kira, Update, PEM, PID test plan IMG_20180208_171423.jpgIMG_20180208_172107.jpgIMG_20180208_171853.jpgIMG_20180208_171932.jpg
                                  Reply  Fri Feb 9 15:29:32 2018, Kira, Update, PEM, PID test plan IMG_20180209_152158.jpg
                               Reply  Thu Feb 15 16:03:57 2018, Kira, Update, PEM, PID test plan 
                               Reply  Thu Feb 22 16:16:43 2018, Kira, Update, PEM, temp sensor input IMG_20180222_160602.jpg
                                  Reply  Mon Feb 26 16:22:10 2018, Kira, Update, PEM, temp sensor input 1.jpgIMG_20180226_154649.jpgIMG_20180226_154720.jpg
                                     Reply  Wed Feb 28 12:31:28 2018, Kira, Update, PEM, temp sensor input 
                                        Reply  Fri Mar 9 12:59:53 2018, Kira, Update, PEM, ADC noise measurement IMG_20180309_114345.jpgIMG_20180309_125153.jpg
                               Reply  Mon Mar 26 16:10:33 2018, Kira, Update, PEM, final setup sketch heater_1_new.pngheater_2_new.pngHeaterCircuit.pdftemp_sensor.png
                                  Reply  Wed Apr 18 12:18:39 2018, Kira, Update, PEM, final setup sketch 7x
                                     Reply  Wed Apr 18 16:59:35 2018, rana, Update, PEM, final setup sketch: EX Seis Screen_Shot_2018-04-18_at_5.20.53_PM.png
                                        Reply  Thu Apr 19 11:29:11 2018, rana, Update, PEM, PID tune Screen_Shot_2018-04-19_at_11.27.08_AM.png
                                           Reply  Mon Apr 23 20:06:35 2018, rana, Update, PEM, PID tune HeaterTest.png
                                              Reply  Thu Apr 26 19:46:26 2018, rana, Update, PEM, PID Quixote dv.pdf
                                                 Reply  Tue May 1 11:15:19 2018, Kira, Update, PEM, PID Quixote IMG_20180501_154826.jpg
                                                    Reply  Tue May 1 15:23:18 2018, Kira, Update, PEM, new ADC channel setup issue IMG_20180501_152014.jpg
                                        Reply  Thu Apr 19 12:23:30 2018, Kira, Update, PEM, final setup sketch update heater_1_new.pngheater_2_new.png
                                           Reply  Thu Apr 19 18:23:51 2018, Kira, Update, PEM, final setup sketch update 
                                              Reply  Fri Apr 20 15:07:45 2018, Kira, Update, PEM, final setup sketch update 
                                           Reply  Thu May 24 16:08:12 2018, Kira, Update, PEM, parts list front_panels.zip
                                              Reply  Tue May 29 16:22:43 2018, Kira, Update, PEM, parts list parts_list_-_Sheet1.pdf1U-panel-2.fpdEX-can-panel.fpd
                                                 Reply  Wed Jun 6 15:12:29 2018, Kira, Update, PEM, parts from lab IMG_20180606_141149.jpgIMG_20180606_141851.jpgIMG_20180606_141912.jpg
                                                    Reply  Mon Jun 18 11:12:23 2018, Kira, Summary, PEM, finishing up work at the lab IMG_20180618_101640.jpgIMG_20180618_093920.jpgIMG_20180618_093932.jpg
                                     Reply  Tue Apr 24 09:10:20 2018, Kira, Update, PEM, final setup sketch EX-can-panel_1.pdf
                                        Reply  Mon Apr 30 15:36:18 2018, Kira, Update, PEM, final setup sketch mount.png
                                           Reply  Tue May 8 10:24:10 2018, Kira, Summary, PEM, plan for this week 
                                              Reply  Fri May 11 11:47:33 2018, johannes, Summary, PEM, Acromag issues 
                                                 Reply  Tue May 15 15:13:23 2018, Kira, Summary, PEM, Acromag issues 
                         Reply  Thu Mar 15 17:33:56 2018, Kira, Update, PEM, test setup IMG_20180315_172512.jpg
                            Reply  Tue Mar 20 16:56:01 2018, Kira, Update, PEM, test setup IMG_20180320_154516.jpgIMG_20180320_145957.jpg
                               Reply  Tue Mar 20 19:48:10 2018, gautam, Update, PEM, test setup 
                               Reply  Fri Mar 23 12:00:20 2018, rana, Update, PEM, test setup 
                               Reply  Fri Mar 23 12:45:08 2018, Kira, Update, PEM, test setup seis_fit.png
                            Reply  Thu May 17 15:38:19 2018, Kira, Update, PEM, test setup with seismometer IMG_20180517_144420.jpgIMG_20180517_145754.jpgIMG_20180517_151956.jpgIMG_20180517_145121.jpg
                               Reply  Fri May 18 14:33:34 2018, Kira, Update, PEM, test setup with seismometer seis_temp_can_on.png
                                  Reply  Mon May 21 14:17:28 2018, Kira, Update, PEM, test setup with seismometer seis-temp.png
                                     Reply  Tue May 22 14:49:03 2018, Kira, Update, PEM, test setup with seismometer seis-temp-2.png
                                        Reply  Wed May 23 14:50:33 2018, Kira, Update, PEM, test setup with seismometer seis-temp-3.pngseis-temp-full-2.png
                                           Reply  Thu May 24 15:18:12 2018, Kira, Update, PEM, PID loop restarted 
                                           Reply  Thu May 31 15:48:16 2018, Kira, Update, PEM, running PID script with seismometer can-on-pid.png
                               Reply  Mon May 21 10:15:35 2018, gautam, Update, PEM, test setup with seismometer 
    Reply  Wed Feb 28 21:14:34 2018, gautam, Summary, PEM, Channel admin tempSensData.png
       Reply  Tue Mar 27 11:11:16 2018, Kira, Update, PEM, Channel setup 
          Reply  Thu Mar 29 17:14:42 2018, Kira, Update, PEM, PID test PID_test.png
             Reply  Fri Mar 30 06:16:45 2018, rana, Update, PEM, PID test 
                Reply  Fri Mar 30 16:10:46 2018, Kira, Update, PEM, PID test 
                   Reply  Wed Apr 4 16:23:10 2018, Kira, Update, PEM, PID test step_response.pngcapacitor.jpgIMG_20180412_120427.jpg
                      Reply  Fri Apr 6 16:17:20 2018, Kira, Update, PEM, PID tuning PID_tuning_progress.png
                         Reply  Fri Apr 6 18:28:57 2018, rana, Update, PEM, PID tuning HeaterTest.png
          Reply  Mon Apr 9 16:30:21 2018, Kira, Update, PEM, MEDM setup MEDM_screen.png
             Reply  Tue Apr 10 15:42:08 2018, Kira, Update, PEM, MEDM setup MEDM_2.png
                Reply  Thu Apr 12 10:15:33 2018, Kira, Update, PEM, MEDM setup MEDM_3.png
                   Reply  Fri Apr 13 00:20:46 2018, rana, Update, PEM, MEDM setup 
Message ID: 13582     Entry time: Thu Jan 25 12:41:18 2018     In reply to: 13575     Reply to this: 13584
Author: Kira 
Type: Update 
Category: PEM 
Subject: Seismometer can insulation test 

We started the actual heating test today and it seems to be working so far. Hoping to heat it to about 40C. We also set up another temperature sensor to measure the lab temperature and connected it to J7, bottom.

Quote:

Gautam and I set up the insulated seismometer can in the lab today. I had previously wired up the two heaters I placed onto the sides of the can in parallel to get a total resistance of 12.5 ohms and then I wrapped the whole can in 3 layers of insulation (k-factor 0.25). We placed it on a large sheet of insulation as to not crush the wires leading out the bottom of the can. I stuck on one of my AD590 sensors to the inside of the can onto the copper lining using duct tape, though this is only a temporary solution. In the future, it would be nice to have some sort of thermal clamp to secure the sensor to the can. To provide power to the heater circuit board and the temperature sensor board, we got a powerstrip and plugged in two power supplies and a function generator into it. The heater circuit (attachment 3) is powered by one of the power supplies and the function generator, while the temperature sensor (attachment 5) is stuck to the side of the can and is powered by the second power supply. The heater circuit's MOSFET (IRF640, attachment 4) is placed on a metal block and sandwiched between two more to make sure it doesn't move around. The temperature sensor is connected by a long BNC cable to the channels in attachment 6.


gautam: we plugged the BNC output of Kira's temperature sensor circuit to J7 on the AA input chassis in 1X2 - this corresponds to ADC1 input 12 in c1ioo. I then made a "PEM" namespace block inside the c1als model, and placed a single CDS filter module inside it (this can be used for calibration purposes). The filter module is named "C1:PEM-SEIS_EX_TEMP", and has the usual CDSfilt channels available. I DQ'ed the output of the filter module (@256 Hz, probably too high, but I'm holding off on a recompile for now). Recompilation and model restart of c1als went smoothly. 

2 bench power supplies are being used for this test, we can think of a more permanent solution later.

**25 Jan noon: Added another filter module, "C1:PEM-SEIS_EX_TEMP", to which Kira is hooking up a second temperature sensor, which will serve as a monitor of the "Ambient" lab temperature. Added DQ channel for the output of this filter module, fixed sampling to 32Hz. Compile and restart went smooth.

 

ELOG V3.1.3-